### v1

# Will Neural Scaling Laws Activate Jevons’ Paradox in AI Labor Markets? A Time-Varying Elasticity of Substitution (VES) Analysis

Abstract
AI industry leaders often use the term “Jevons’ Paradox.” We explore the significance of this term for artificial intelligence adoption through a time-varying elasticity of substitution framework. We develop a model connecting AI development to labor substitution through four key mechanisms: (1) increased effective computational capacity from both hardware and algorithmic improvements; (2) AI capabilities that rise logarithmically with computation following established neural scaling laws; (3) declining marginal computational costs leading to lower AI prices through competitive pressure; and (4) a resulting increase in the elasticity of substitution between AI and human labor over time. Our time-varying elasticity of substitution (VES) framework, incorporating the Gørtz identity, yields analytical conditions for market transformation dynamics. This work provides a simple framework to help assess the economic reasoning behind industry claims that AI will increasingly substitute for human labor across diverse economic sectors.
1 Introduction
Many industry leaders have appealed to Jevons’ Paradox (1866) where price declines lead to more usage when talking about the progress of AI. Also, industry leaders have stated that the AI scaling laws will lead to quality improvements (Kaplan et al., 2020; Hoffmann et al., 2022) that will also lead to more usage. [^1] For example, Sam Altman (2025), in his blog entitled “Three Observations,” stated that “1. The intelligence of an AI model roughly equals the log of the resources used to train and run it. 2. The cost to use a given level of AI falls about 10X every 12 months, and lower prices lead to much more use. 3. The socioeconomic value of linearly increasing intelligence is super-exponential in nature.” The first point restates the AI scaling law while the second point is consistent with Jevons’ Paradox. Satya Nadella (2025) stated that “Jevons paradox strikes again! As AI gets more efficient and accessible, we will see its use skyrocket, turning it into a commodity we just can’t get enough of.” Driving much of this has been the decreased marginal cost of computation and competition that links the lower marginal costs to the lower cost of AI. [^2] Ho et al. (2024) report that the “… compute required to reach a set performance threshold has halved approximately every 8 months, … .” Is Jevons’ Paradox mere industry jargon, or does it provide economically meaningful insights relevant to AI progress?
Jevons’ Paradox occurs when increased efficiency in a resource leads to greater total demand due to both intensive margins (existing users consuming more) and extensive margins (new users or applications entering the market which may take time). For example, the extensive margin of uses from electricity has dwarfed the initial uses of electricity for lighting, but many uses took many years to emerge. In elasticity terms, this implies a long-term price elasticity of demand less than $-1$ (i.e., highly elastic, $E<-1$), where the percentage increase in quantity demanded exceeds the percentage decrease in price. The AI context involves both pecuniary price decreases as well as quality increases for a reduction in the quality-adjusted price.
This analysis develops a framework that incorporates these statements so as to make the underlying logic clearer. Specifically, we examine market transitions to AI-produced goods using an elasticity of substitution framework, drawing on the identity presented in Gørtz (1977). [^3] Smith (1979) points out that the identity was not necessarily new, but implied by Hicks and Allen (1934), although the derivation was new. We demonstrate conditions under which complete market transformation occurs, highlighting the role of both price elasticities and substitution elasticities.
To address the question of Jevons’ Paradox in the context of AI competing with human labor, and whether the quantity demanded of an AI-provided good rises relatively more than its price decreases, we analyze this using a CES utility function (Arrow et al., 1961) and an identity between price elasticity and the elasticity of substitution from Gørtz (1977).
We further explore the effects of exponential declines in the marginal cost of effective computation which via competition leads to exponential declines in the price of AI-provided goods and services as well as the increased quality of AI goods and services which flows from the increased effective computation created by the lower marginal costs via the AI “scaling laws.” The combination of exponential declines in the price of AI-provided goods and services and increases in quality (increased ability to substitute) leads to AI adoption with an early phase dominated by price making way for a mature phase that uses quality increases for market domination. In fact, unless AI quality rises sufficiently (elasticity of substitution above 1), no level of AI price declines (provided $p_{A}>0$) will result in complete capture of markets.
This paper makes three primary contributions: First, we formalize the connection between elasticity of substitution and own-price elasticities in AI-human substitution using the Gørtz (1977) identity. Second, we derive analytical conditions for market transformation using a variable elasticity of substitution specification. Third, as an aid to categorization, we set out a five-phase adoption process with distinct economic mechanisms governing each phase that highlights when Jevons’ Paradox occurs.
2 Jevons’ Paradox and Elasticity
Assume a CES utility function for consumers choosing between human-produced (H) and AI-produced (A) goods in (1),
$U$ $=(H^{\rho}+A^{\rho})^{1/\rho}$ (1)
where $\rho=\frac{\sigma-1}{\sigma}$ and $\sigma$ represents the elasticity of substitution. Gørtz (1977) presented an identity in (2) that relates the own price elasticities and the elasticity of substitution,
$\sigma+E_{AA}+E_{HH}=-1$ (2)
where for this problem $E_{AA}$ is the own-price elasticity of AI goods, and $E_{HH}$ is the own-price elasticity of human goods. The Gørtz identity provides an analytical link between the elasticity of substitution and own-price elasticities in this AI-human substitution framework. This relation allows characterization of market transitions more completely than examining either elasticity measure in isolation. For AI-produced goods, the elasticity identity reveals boundary conditions for market dominance that depend on both the intrinsic substitutability between AI and human production and the market’s response to price changes for each good, service, or task.
Given normalized human prices ($p_{H}=1$) and AI prices $p_{A}\in(0,1)$, the AI expenditure or revenue share, $r_{A}$, appears in (3). [^4] The ratio of the AI price to human prices is not that easy to quantify. First, AI prices involve the cost of the tokens for input and output. Second, queries differ across numbers of tokens required. Third, many queries also involve a human and so the AI cost is really a hybrid AI-human cost. Fourth, many tasks may require multiple queries. Finally, the cost of a human query also varies by task. Therefore, the ratio of the AI price to the human price is not obvious in magnitude, but should be less than 1.
$r_{A}$ $=\frac{1}{1+p_{A}^{\,\sigma-1}}$ (3)
What happens as the price of AI declines? Assume $p_{A}\to 0$,
$r_{A}$ $\to 1,\text{ for }\sigma>1$ (4)
This reveals that no matter what the price of AI, if AI does not have sufficient quality to substitute for human products, services, or tasks, it cannot achieve market domination in expenditures $r_{A}$.
It may aid understanding of the role of price and the elasticity of substitution to consider the elasticity of the revenue share $r_{A}$ to the AI price $p_{A}$ in (5).
$E_{r_{A},p_{A}}=-\frac{(\sigma-1)p_{A}^{\sigma-1}}{1+p_{A}^{\sigma-1}}$ (5)
Clearly, the revenue rises as price decreases when $\sigma>1$. However, for $\sigma\in(1,2)$, a concave relation exists for $p_{A}^{\sigma-1}$ and this becomes convex for $\sigma>2$. This arises because for $p_{A}<1$, $\sigma\in(1,2)$ moves $p_{A}^{\sigma-1}$ towards 1, while $\sigma>2$, moves $p_{A}^{\sigma-1}$ towards 0.
For the revenue share $r_{A}$ this aligns with Gørtz’s identity since one can substitute out $E_{HH}$ to arrive at (6).
$E_{AA}$ $=-\sigma+(\sigma-1)r_{A}$ (6)
Equation (6) shows that as the AI market share approaches complete dominance ($r_{A}\to 1$), the own-price elasticity converges to unity ($E_{AA}\to-1$), representing the cessation of Jevons’ Paradox. This transition from elastic to unitary elastic demand signifies market saturation, where further price reductions no longer generate proportionally larger quantity increases. The convergence occurs precisely when substitution opportunities are exhausted, representing a natural boundary condition in the AI adoption process.
3 Exponential Declines in the AI Price, Scaling Laws, and Increased Elasticity of Substitution (VES)
This analysis examines AI market penetration dynamics under the further assumption of exponential growth in computational capacity and exponential decline in AI prices. We derive closed-form solutions that demonstrate the interaction between computational growth rates and price dynamics in determining market transitions.
Assume computational capacity grows exponentially as in (7),
$\ln\left(\frac{C(t)}{C(0)}\right)$ $=g\cdot t$ (7)
where $g$ is the computational growth rate. As mentioned earlier, Ho et al. (2024) document that the compute required to reach a given level of performance for large language models declines by 50% every 8 months. This translates to a gain in effective computational capacity of $1.04$ in terms of continuous compounding per year. Accordingly, one could entertain a prior around $g\approx 1$. This ignores the substantial gains from better hardware as well as the large amounts of investment in AI. Against this background, the “scaling laws” such as Kaplan et al. (2020) and Hoffmann et al. (2022) indicate that AI capabilities rise logarithmically with computation. In these works computation is proportional to the number of parameters times the data and thus computation implies scaling of these.
Although most economic models rely on constant elasticity of substitution (CES) models to capture trade-offs among choices, a CES approach does not naturally capture the increasing capabilities of AI systems over time. The use of a variable elasticity of substitution (VES) framework addresses this (Revankar, 1971).
Specifically, assume that improvements in AI capabilities increase the elasticity of substitution by means of the scaling laws which show a linear improvement in AI performance as inputs (in this case computation) increase logarithmically as in (8). Note, the actual value of $\delta>0$ could differ across goods, services, tasks, contexts, or organizations.
$\sigma(t)$ $=\delta\cdot\ln\left(\frac{C(t)}{C(0)}\right)$ (8)
Due to the exponential growth in computation in (7), (8) reduces to (9).
$\sigma(t)$ $=(\delta\cdot g)\cdot t=\phi\cdot t,\text{ where }\phi=\delta\cdot g$ (9)
This yields an elasticity of substitution in (9) where the constant $\phi=\delta\cdot g$ represents the compound effect of elasticity of substitution gain through greater computation times the rate of additional computation. [^5] As long as the elasticity of substitution rises with computation (scaling laws) and computation rises over time (Moore’s Law for hardware, Erdil and Besiroglu (2022) for software, Ho et al. (2024) for language models), the exact form of the relation does not affect the outcome, only the timing. A linear or concave relation seems conservative. However, if we rely on Sam Altman’s assertion (point 3) that “The socioeconomic value of linearly increasing intelligence is super-exponential in nature.”, $\sigma(t)$ could even rise at a high rate.
The compound effect parameter $\phi$ captures two forces driving AI adoption. Namely, the rate at which computation increases ($g$) and the sensitivity of the elasticity of substitution $\delta$ to computational improvements. In other words, $\phi$ specifies the elasticity of substitution growth rate. This form links the technical AI scaling laws literature with economic substitution theory. Note, the models improve in terms of loss with repect to inputs at different reported rates, Kaplan et al. (2020) reported a rate of loss decline of around $0.05$. So if this directly translated into $\delta$, one could entertain a prior on $\delta$ of around $0.05$. However, as mentioned above, this likely varies substantially across various uses. [^6] We use the word “prior” in conjunction with the parameters $g$, $\delta$ to convey a sense of uncertainty as to their values. The model treats the parameters as constants. One could certainly place an informative prior on these parameters and simulate across the potential outcomes which might yield a more dispersed, non-symmetric distribution. We do not pursue this further as the goal is create a simple framework while qualitatively acknowledging uncertainty.
Consistent with Ho et al. (2024), assume exponentially declining AI prices in (10),
$p_{A}(t)$ $=p_{A}(0)\cdot\exp(-d\cdot t)$ (10)
where $10<p_{A}(0)<$ 1 represents the initial price ratio between AI and human goods. [^7] Although we model the declining marginal cost cost of computation and the increased quantity of computation separately, they obviously are related through the demand equation. However, there are other shifters of demand and so we will retain the simplicity of separate equations. In addition, to the degree that AI work, especially when it has a low elasticity of substitution for human work, in reality is hybrid AI-human work. Therefore, pecuniary AI price declines may overstate the price declines in actual tasks.
Substituting this modified price function into equation (3) while maintaining the variable elasticity of substitution from equation (8), leads to a generalized logistic function (11) that represents the revenue share $r_{A}(t)$ based on parameter values for $a,b,c$ in (12).
$r_{A}(t)$ $=\frac{1}{1+\exp\Bigl{(}a+b\cdot t+c\cdot t^{2}\Bigr{)}}$ (11)
$a$ $=-\ln\bigl{(}p_{A}(0)\bigr{)},\quad b=d+\phi\cdot\ln\bigl{(}p_{A}(0)\bigr{)},\quad c=-d\cdot\phi$ (12)
The revenue share $r_{A}(t)$ will monotonically increase over time after time $t_{0}$ (where $t_{0}$ could be 0, 1, or some other value) provided the initial price is low enough as shown in (13).
$p_{A}(0)<\exp\left(d\cdot\left(2\cdot t_{0}-\frac{1}{\phi}\right)\right)$ (13)
This comes from taking the derivative of the argument in the exponential in (11) ($b+2\cdot c\cdot t$) and making sure that it is strictly negative for $t\geq t_{0}$.
Based on the model, as an aid to categorization, AI adoption could be thought of as occurring in five phases as presented below:
  1. 1.
    Phase 1 represents the complementarity-dominated regime where $\sigma(t)<1$ since $p_{A}<p_{H}=1$. AI enhances rather than replaces human production, with adoption driven primarily by initial price advantage. Specifically, when $t$ is near 0 market penetration is driven by both price effects and the initial low market share. In other words, AI firms must try to begin with a low price to initiate or jumpstart the market. During this phase, perhaps a long phase, AI goods, services, or tasks are complements to the human counterparts. Even if $p_{A}(t)$ is almost 0, human goods, services, or tasks will have positive share or $r_{A}<0.5$. Jevons’ Paradox does not occur during this phase. Currently, tasks such as coding display relatively large shares of AI, but under 50%. [^8] In the October 29th 2024 earnings call, Google CEO Sundar Pichai stated “We’re also using AI internally to improve our coding processes, which is boosting productivity and efficiency. Today, more than a quarter of all new code at Google is generated by AI, then reviewed and accepted by engineers.” In Patel (2025), Nadella states his beliefs in human-AI complementarity, but Nadella (2025) also promotes Jevons’ Paradox which indicates they are substitutes. Although both could be true, given factors outside of the simple model here via large-scale economic growth. Altman (2025) often suggests AI will exit this first phase.
  2. 2.
    Phase 2 marks the critical transition point where $\sigma(t)=1$, corresponding to unitary elasticity of substitution. This occurs when $t^{*}=\phi^{-1}$. This marks the tipping point where the elasticity of substitution between AI and human goods, services, or tasks is equal so that $\sigma(t)=1$. This phase exhibits properties analogous to Cobb-Douglas production, with constant quantity and revenue shares across AI and human inputs despite changing relative prices. In fact, this corresponds to $r_{A}=0.5$, a large market share. At this point, Jevons’ Paradox is at the cusp of occurring.
  3. 3.
    Phase 3 initiates true Jevons’ dynamics in revenue share where $\sigma(t)>1$ happens when $t>t^{*}$). The price elasticity satisfies $E_{AA}<-1$, indicating that percentage increases in quantity exceed percentage decreases in price. Jevons’ Paradox comes into play and price decreases actually increase revenue by spurring more usage. As $t$ increases, the term $-d\cdot\phi\cdot t^{2}$ dominates, representing the increasing influence of computational scaling and lower prices on market transformation. In this phase the AI market share is between 50 and 100% ($10.5<r_{A}<$ 1).
  4. 4.
    Phase 4 begins when $\sigma(t)\geq 2$ at $t\geq 2\cdot t^{*}$ and thus accentuates Jevons’ dynamics in revenue share. In the beginning of this phase, when $\sigma=2$, the revenue share shows a simple relation in (14) with $p_{A}$, the price of AI, based on substitution of $\sigma=2$ into (3). Note, if $p_{A}$ is already low by the time $\sigma=2$, the revenue share $r_{A}$ may already risen close to 1. The converse of (14) appears in (15) where the price $p_{A}$ shows a simple relation with revenue share, $r_{A}$. Also, the elasticity of revenue share to price $E_{r_{A},p_{A}}$ has a simple form in (16), based on substitution of $\sigma=2$ into (5). For low values of $p_{A}$, the elasticity may have already have low magnitude as lowering the price more will not bring forth more revenue.
    $r_{A}$ $=\dfrac{1}{1+p_{A}}$ (14)
    $p_{A}$ $=\frac{1-r_{A}}{r_{A}}$ (15)
    $E_{r_{A},p_{A}}$ $=-\dfrac{p_{A}}{{1+p_{A}}}=-p_{A}\cdot r_{A}$ (16)
    Of course, as time progresses and the elasticity of substitution $\sigma(t)$ increases beyond 2 while the price $p_{A}$ continues to decrease, the revenue share will continue its ascent.
  5. 5.
    Phase 5 represents market saturation within the constrained partial equilibrium framework, where $r_{A}=1$. At this boundary, Jevons’ Paradox no longer applies as the price elasticity equals $-1$, indicating unitary elasticity. Of course, if AI spurs economic growth the overall market could expand and Jevons’ Paradox could continue to hold. This sentiment has been put forth by industry leaders, but addressing this lies outside the scope of this simple model.
4 Conclusion
Frequent invocation of Jevons’ Paradox by AI leaders may appear as typical industry jargon devoid of much meaning, but as shown here it equates to stating that AI can substantially or completely substitute for human provision of goods, services, or tasks. Given the importance of Jevons’ Paradox to the progress of AI, it seems timely to examine it in more detail.
This analysis provides a framework that incorporates these statements to make the underlying logic clearer. Specifically, we examine market transitions to AI-produced goods using an elasticity of substitution framework, drawing on the identity presented in Gørtz (1977). We introduce a variable elasticity of substitution (VES) model where the increased performance of AI based on the scaling laws expands as a function of computational growth that leads to the elasticity of substitution $\sigma(t)$ growing over time. The growth in $\sigma(t)$ coupled with falling prices for AI driven by competition and the lower marginal costs of computation (stemming from hardware, algorithmic, and software progress) create an environment where AI becomes better and more inexpensive relative to humans over time.
As a categorization of AI progress, we set forth a five-phase model where (1) initial adoption depends critically on sufficiently low starting prices relative to substitution sensitivity; (2) transition points between phases are determined by parameters linking computational scaling to substitution elasticity; (3) Jevons’ Paradox occurs in terms of increases in revenue with price decreases when $\sigma(t)>1$; (4) A stronger form of Jevons’ Paradox occurs when $\sigma(t)>2$; and (5) market saturation eventually leads to unitary price elasticity (no Jevons’ Paradox).
Having presented this model to understand the perspective of AI proponents, many alternative perspectives exist as to the reasons why explosive AI growth may or may not occur (Erdil and Besiroglu, 2024). The simple partial equilibrium model presented above does not account for the many positive or negative feedback loops in a more general equilibrium that could accelerate or impede adoption. For example, we held the price of the human product at 1, regardless of the context, whereas this may change in response to AI competition. We also do not discuss the differential changes in substitution and associated feedback loops that could occur across products, services, or tasks as AI, through the mechanism of high-quality data availability (e.g., programming, science). In addition, the optimistic view of AI stresses the increase in overall wealth and quality-of-life potential which could feedback into choices. Finally, if AI grows income and wealth, usage could continue to grow despite having captured the entire market which could lead to a continuation of Jevons’ Paradox. In addition, we do not discuss possible impediments to AI development such as declining quantity and quality of data, energy usage, capital costs, scaling laws failing to hold, frictions, and so forth.
Despite the barriers that AI adoption could face, investor revealed preference indicates their latent, positive AI forecast. For example, Apple recently announced \$500B of AI related investments (Tobin and Karaian, 2025), and many of the leading firms have set forth \$300B in planned investments (Waters and Bradshaw, 2025). Supporting their optimism, OpenAI now has over 400 million weekly active users (Mishkin and Nellis, 2025).
The influence of Jevons (1866), research on scaling laws (Kaplan et al. 2020; Hoffmann et al., 2022), which at the time was “academic,” established concepts of marginal cost as well as competition, and Moore’s Law (Moore, 1965) may have motivated these large investments which cumulatively exceed \$1T. Whether this leads to the greatest investment bubble of all time or an age of “super abundance” remains to be seen as the scope of the opportunity may increase the scope for possible error. Given the motivation for these investments often invokes well-established economic ideas and technical research, a quote from Keynes (1936) may prove apt.
“The ideas of economists and political philosophers, both when they are right and when they are wrong, are more powerful than is commonly understood. Indeed the world is ruled by little else. Practical men, who believe themselves to be quite exempt from any intellectual influences, are usually the slaves of some defunct economist. Madmen in authority, who hear voices in the air, are distilling their frenzy from some academic scribbler of a few years back. I am sure that the power of vested interests is vastly exaggerated compared with the gradual encroachment of ideas.”